OpenStudy (anonymous):

dy/dx = (cos x)e^(sin x), y(0) = 0, solve the initial value problem

OpenStudy (freckles):

so you to integrate both sides to find y

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{dy}{dx} dx=\int\limits_{}^{}\cos(x)e^{\sin(x)} dx\]

OpenStudy (freckles):

use a sub on the right hand side

OpenStudy (freckles):

are you there?

OpenStudy (anonymous):

yes sorry. having computer problems

OpenStudy (anonymous):

u sub. would it be u=sinx du=cosx?

OpenStudy (freckles):

well du=cos(x)dx

OpenStudy (freckles):

and yes

OpenStudy (freckles):

\[y=\int\limits_{}^{}e^{u} du\]

OpenStudy (anonymous):

answer being e^sin (x) -1 when you plug in initial conditin of y(0) =0 yes?

OpenStudy (freckles):

looks good

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

would dy/dx = x/y be

OpenStudy (anonymous):

\[\int\limits y dy = \int\limits x dx\]